A zoo has a circular pool for its seals. The diameter of the pool is 32 feet.How much fence is needed to enclose the pool?
1
Expert's answer
2016-09-08T08:13:03-0400
2*Pi*R=D*Pi=32*Pi=32*3.1415926=100.53 feet of fence is needed. Here R is the radius of the circle, D is the diameter of the circle, Pi=3.1415926...